OpenStudy (anonymous):

aDivide the following polynomial using synthetic division, then place the answer in the proper location on the grid. Write answer in descending powers of x. (x^3 - 7x^2 + 8 ) ÷ (x + 1)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Let f(x) = x^3-7x^2+7x+15 By inspection (ie: guessing): f(-1) = -1 - 7 - 7 + 15 =0 By the factor theorem is f(-1) = 0 then (x+1) is a factor of f(x). x^3-7x^2+7x+15 = (x+1)(Ax^2+Bx+C) where A, B and C are constants to be found either by long division or by comparing coefficients. I will be doing the latter as it's easier to show. Coefficients of x^3: 1 = A Coefficients of x^0 (constant term): 15 = C Coefficients of x^2: -7 = A + B , B = -8 Thus f(x) = (x+1)(x^2-8x+15) We need to now see if that quadratic factors and it will if the discriminant is a complete square: b^2-4ac = 8^2 - 4*1*15 = 64-60 = 4 so this does factor: and so the factored expression is f(x) = (x+1)(x-3)(x-5)

OpenStudy (anonymous):

i hope that helps u @gary4

OpenStudy (precal):

here is your set up for synthetic division